# Vision Beyond Boundaries: An Initial Design Space of Domain-specific Large Vision Models in Human-robot Interaction

### v3

@@ Abstract.
The emergence of large vision models (LVMs) is following in the footsteps of the recent prosperity of Large Language Models (LLMs) in following years. However, there’s a noticeable gap in structured research applying LVMs to human-robot interaction (HRI), despite extensive evidence supporting the efficacy of vision models in enhancing interactions between humans and robots. Recognizing the vast and anticipated potential, we introduce an initial design space that incorporates domain-specific LVMs, chosen for their superior performance over normal models. We delve into three primary dimensions: HRI contexts, vision-based tasks, and specific domains. The empirical evaluation was implemented among 15 experts across six evaluated metrics, showcasing the primary efficacy in relevant decision-making scenarios. We explore the process of ideation and potential application scenarios, envisioning this design space as a foundational guideline for future HRI system design, emphasizing accurate domain alignment and model selection.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Vision for roboticsccs: Computing methodologies Computer visionccs: Human-centered computing HCI theory, concepts and models
@@ 3.3. Specific domains
  • Healthcare: HRI is revolutionizing health such as robotic-assisted surgery, where precision and reduced invasiveness are critical. Robots like exoskeletons are also being deployed for rehabilitation, providing physical therapy with adaptive routines. The superiority of LVMs in detecting anomalies in medication-related imaging and forecasting disease progression will enhancing the decision-making.
  • Automotive: LVMs are central in advancing the autonomous vehicles. The real-time object and pedestrian detection, will enable safer navigation and interaction with the environment. Furthermore, LVMs enhance the design of ergonomic vehicle interiors, thereby elevating user experience and safety in driver-assist systems.
  • Manufacturing: LVMs enable robots to execute precise quality inspections, detect defects, and ensure accurate assembly, collaborating with human operators. In addition, the process monitoring and maintenance significantly improve the efficiency and workplace safety.
  • Entertainment: LVMs in HRI are anticipated to vastly improve entertainment experiences through their prowess in generating visual content, including enhanced video and gaming experiences. Moreover, the capacity for analyzing visual data, like predicting human intentions, is set to further enrich the quality of interactive experiences between humans and robots.
  • Security: In critical situations, robots are enabled to perform continuous surveillance, recognize suspicious activities, and identify potential threats with desirable accuracy. Also, utilizing advanced facial recognition and behavior analysis from LVMs to ensure human safety and security publicly.
  • Agriculture: LVMs enable agricultural robots (such as drones) to accurately identify and classify crops, assess plant health, predict incoming yields, and detect pests or diseases, optimizing the cultivating and harvesting.
  • Education: This domain has been ubiquitously probed in HRI. With the aid of LVMs, educational robots are to precisely recognize and respond to manifold gestures and expressions from humans, facilitating more engaging and personalized teaching performance. Some vision-related techniques such as virtual reality/augmented reality are able to make learning more immersive and accessible.
  • Social Interaction: Social robotics is a crucial facet of HRI. In social settings, LVMs enable robots to interpret and respond to human emotions and social cues, fostering more natural and empathetic communication and engagement. A wide-used case is eldercare, where the social robots can provide companionship and emotional support, tailored to individual needs.
@@ 4.1. Study Design
Exploring the design space’s dimensions provides valuable insights into LVM diversity for HRI researchers and practitioners. However, it doesn’t reveal how these variations affect individual responses. To address this, we conducted an expert evaluation to gather empirical evidence for developing HRI systems that evoke specific reactions to LVMs. We engaged 15 participants aged 24 to 42 (7 male, 6 female, 2 prefer not to say, M = 30.33, SD = 4.66) through email invitations and personal outreach. Participants had several years of experience in relevant areas (see Figure  2.A), enhancing the study’s reliability. After a brief introduction to the design space, we used a questionnaire to explore experts’ perspectives on its three dimensions and the design space itself. This study employed a questionnaire-based evaluation and quantitative analysis to measure five perceived metrics from existing literature (Gries, 2004; Kalegina et al., 2018; Baartmans et al., 2020; Haggerty et al., 2011). The questionnaire included demographic information, an overview of the design space (see Figure  1), and targeted questions to gauge perceptions of likeability, trustworthiness, usefulness, intent to use, and comprehensiveness, using a 7-point Likert scale. The questions are detailed in the appendix.
@@ Acknowledgements.
Appendix A The questions designed for the study
  • How much do you like this dimension/design space? (Very dislike – Very like)
  • How trustworthy do you consider this dimension/design space to be? (Very untrustworthy – Very trustworthy)
  • How useful do you find this dimension/design space in guiding or inspiring future research in related areas? (Very useless – Very useful)
  • How likely are you to utilize this dimension/design space in designing future HRI systems that incorporate LVMs? (Very unlikely – Very likely)
  • How comprehensive do you find the dimension/design space in covering the aspects relevant to the future research in related areas? (Very incomprehensive – Very comprehensive)

### v4

@@ Abstract.
The emergence of large vision models (LVMs) is following in the footsteps of the recent prosperity of Large Language Models (LLMs) in following years. However, there’s a noticeable gap in structured research applying LVMs to human-robot interaction (HRI), despite extensive evidence supporting the efficacy of vision models in enhancing interactions between humans and robots. Recognizing the vast and anticipated potential, we introduce an initial design space that incorporates domain-specific LVMs, chosen for their superior performance over normal models. We delve into three primary dimensions: HRI contexts, vision-based tasks, and specific domains. The empirical evaluation was implemented among 15 experts across five evaluated metrics, showcasing the primary efficacy in relevant decision-making scenarios. We explore the process of ideation and potential application scenarios, envisioning this design space as a foundational guideline for future HRI system design, emphasizing accurate domain alignment and model selection.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 20XX; XXXXXXX, XXbooktitle: Conference ’18: ACM Conference on Some Themes, June 03–05, 20XX, XXXXXXXX, XXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Vision for roboticsccs: Computing methodologies Computer visionccs: Human-centered computing HCI theory, concepts and models
@@ 3.3. Specific domains
  • Healthcare: HRI is revolutionizing health such as robotic-assisted surgery, where precision and reduced invasiveness are critical (Lee et al., 2012; Esterwood and Robert, 2020). Robots like exoskeletons are also being deployed for rehabilitation, providing physical therapy with adaptive routines. The superiority of LVMs in detecting anomalies in medication-related imaging and forecasting disease progression will enhancing the decision-making.
  • Automotive: LVMs are central in advancing the autonomous vehicles (Heydaryan et al., 2018; Basu and Singhal, 2016). The real-time object and pedestrian detection, will enable safer navigation and interaction with the environment. Furthermore, LVMs enhance the design of ergonomic vehicle interiors, thereby elevating user experience and safety in driver-assist systems.
  • Manufacturing: LVMs enable robots to execute precise quality inspections, detect defects, and ensure accurate assembly, collaborating with human operators (Marvel et al., 2020; Sadrfaridpour and Wang, 2017; Tan et al., 2009). In addition, the process monitoring and maintenance significantly improve the efficiency and workplace safety.
  • Entertainment: LVMs in HRI are anticipated to vastly improve entertainment experiences through their prowess in generating visual content, including enhanced video and gaming experiences (Budiharto et al., 2017; Pollmann et al., 2023). Moreover, the capacity for analyzing visual data, like predicting human intentions, is set to further enrich the quality of interactive experiences between humans and robots.
  • Security: In critical situations, robots are enabled to perform continuous surveillance, recognize suspicious activities, and identify potential threats with desirable accuracy (Agrawal and Williams, 2017; Lopez et al., 2017). Also, utilizing advanced facial recognition and behavior analysis from LVMs to ensure human safety and security publicly.
  • Agriculture: LVMs enable agricultural robots (such as drones) to accurately identify and classify crops, assess plant health, predict incoming yields, and detect pests or diseases, optimizing the cultivating and harvesting (Vasconez et al., 2019; Adamides et al., 2017).
  • Education: This domain has been ubiquitously probed in HRI. With the aid of LVMs, educational robots are to precisely recognize and respond to manifold gestures and expressions from humans, facilitating more engaging and personalized teaching performance (Tekerek, 2009; Cui et al., 2022). Some vision-related techniques such as virtual reality/augmented reality are able to make learning more immersive and accessible.
  • Social Interaction: Social robotics is a crucial facet of HRI (Yan et al., 2014; Kanda and Ishiguro, 2017; Lemaignan et al., 2017). In social settings, LVMs enable robots to interpret and respond to human emotions and social cues, fostering more natural and empathetic communication and engagement. A wide-used case is eldercare, where the social robots can provide companionship and emotional support, tailored to individual needs.
@@ 4.1. Study Design
Exploring the design space’s dimensions provides valuable insights into LVM diversity for HRI researchers and practitioners. However, it doesn’t reveal how these variations affect individual responses. To address this, we conducted an expert evaluation to gather empirical evidence for developing HRI systems that evoke specific reactions to LVMs. We engaged 15 participants aged 24 to 42 (7 male, 6 female, 2 prefer not to say, M = 30.33, SD = 4.66) through email invitations and personal outreach. Participants had several years of experience in relevant areas (see Figure  2.A), enhancing the study’s reliability. After a brief introduction to the design space, we used a questionnaire to explore experts’ perspectives on its three dimensions and the design space itself. This study employed a questionnaire-based evaluation and quantitative analysis to measure five perceived metrics from existing literature (Gries, 2004; Kalegina et al., 2018; Baartmans et al., 2020; Haggerty et al., 2011). The questionnaire included demographic information, an overview of the design space (see Figure  1), and targeted questions to gauge perceptions of likeability, trustworthiness, usefulness, intent to use, and comprehensiveness, using a 7-point Likert scale. The questions are detailed in the supplementary material.